OpenStudy (anonymous):

Which of the following is one of the factors of 3x2 + 45x + 150?

OpenStudy (anonymous):

how do i do this?

OpenStudy (cwrw238):

first take 3 out = 3(x^2 + 15x + 50) = 3 (x+10)(x+5)

OpenStudy (anonymous):

5 All of the above x – 10 x + 5

OpenStudy (anonymous):

so its x+5

OpenStudy (anonymous):

x+5

OpenStudy (anonymous):

3x2+45x+150=3(x+5)(x+10) Answer = x + 5

OpenStudy (cwrw238):

to factor whats in the brackets you need 2 numbers whose product is 50 and whose sum is 15 - thasts 5 and 10

OpenStudy (cwrw238):

yes correct option is x+5